RC Update and Run Plan for March 19, 2020
Current RC: Jay Benesch email@example.com
RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection
- Hall is in restricted access.
- Let the RC, Jay Benesch, and Walter Kellner know about any non-pivot work you'd like to take place during the March 13-20 down in Hall C. (Please CC Brad as well.)
- Field probe died around end of swing shift, halting mapping program. Survey delayed.
Bill Henry and others have been mapping with one plane probe. Jian-Ping drove the 3D to W&M to recalibrate when we couldn't contact Magnet Measurement. He got back about 1800 3/18 and work proceeded all night.
- B and C laser seeds were switched this morning. https://logbooks.jlab.org/entry/3806164
- The work items are currently:
- A 3He compressor checkout for the x > 1 experiment next year (Meekins/RadCon);
RC mentioned this at 0800 meeting 3/18. Dave Hamlette said he'd coordinate with Meekins.
- Replace a gauge in the SHMS detector hut and put a lead blanket around it (Kenyon) DONE
- SOLID detector test the scope of which is defined in the HClist http://devweb.acc.jlab.org/CSUEApps/hclist/task/1809 (Malace et al).
- Vote for your favorite T-shirt design https://misportal.jlab.org/ul/apps/jag/tshirt/index.cfm I prefer Staphany's JLAB word cloud, entry 11, for reasons which will become obvious if you look closely.
Current Run Plan:
- continue Solid prototype installation
- Field and Compass measurement continue until 0900
- pointing Surveys for SHMS, HMS Thursday 0900-1500
- Thursday 1500-1600+ techs crane down optical bench and install staircase to second level
- Install new downstream beamline window spool Thursday, pump down beamlines Friday or Saturday (install protective window guards)
- ESH evaluation of laser platform work 0815 Friday
- Re-install upstream Be window Friday or Saturday with RadCon support.
- Install reference cell, leak check, etc.
- Install target cell and ladder as soon as possible
- Recommission target and start spinning up ASAP (through SWING and OWL as needed)
Notes for the Future
- At d2n startup, it would be a good idea to quantitatively check the dispersion at the target. I (D.Mack) don't think we need a tight spec, but if it's as sloppy as 1 cm/% (hence a decent spectrometer), then we're more at risk for things like how much the beam moves when the locks are off, or scraping from an off-energy tail on the beam. Jay:MQA3C13 can be adjusted to change horizontal dispersion.
- 'Filler' runs for next spin up / target-down period (Don't want to forget about this!):
- DONE 1 hour run with target out for measure of the background just from the beryllium windows. (Needs SHMS, btw.)
- Be sure to restart main alarm handlers and turn up speaker volume before operating magnets! 
- Be sure to restart target alarm handler and turn up speaker volume before operating target! 
- pNMR issues
- Software/driver issue needs to be resolved. Checkout during next non-production down time (Moller or SHMS fix)
- EPR problems
-  saw some progress on Tuesday , but issues remain.
- Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares).
- It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
- Kepco power supply
- PS commissioning/installation (Todd, Bill) (Bill got a filter design from the vendor and has ordered parts.)
- New filter (for 180 Hz) to install for current mode operation fix. Can always go to voltage mode if it doesn't work.
- NMR Calibration Constants
- Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted
-      (Bill, Junhao, Arun)
- Add list of calibration runs and what constants apply to what production runs here: PolHe3_Polarization_Measurements
- Beam trips
- Persistent BLM and ion chamber trips
- General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
- NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
- Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
- USB bridge
- Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
- Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
- Ensure issues on Sat, March 14 have been corrected 
- IOC issues
- Green now: Rebooted 3/2/20 with Dave Gaskell assistance. OPs wished to reboot hung iochc10 , noted in December as well 
- This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.
Opportunistic Access Jobs
None at the moment.
Bluejeans remote connection information
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-126.96.36.199 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920